OpenStudy (anonymous):

Giving medal to best answer? A survey of students in Oklahoma found that 60% of the students have seen a tornado on television. Of those who have seen a tornado on television, 28% have had a tornado in their hometown. What percent of the students surveyed have seen a tornado on television and have had one in their hometown?

OpenStudy (gorv):

let total student = 100

OpenStudy (gorv):

what will be its 60 %??

OpenStudy (anonymous):

40 students?

OpenStudy (anonymous):

If i'm not mistaken it should be 60% = 1/6 x 28% = 14/50 -> 14/300 = 4.6%

OpenStudy (anonymous):

NVM it didn't work

OpenStudy (anonymous):

Oh, I got the answer. But, thankyou anyways! It was 16.8%

OpenStudy (anonymous):

I did the math wrong but the equation was correct haha

OpenStudy (anonymous):

You have your first percentage multiplied by your second percentage

OpenStudy (anonymous):

yes, you can